A, B, and C are on a betting game. B loses P350 of his money to A. As a result, A now has twice as much as what is left with B. Then, C loses P700 to B. As a consequence, C now has only one-third as much money as B would then have. If A loses P210 to C, C will have as much money as A would have left. How much did each have at the start?


1
Expert's answer
2022-04-07T17:17:53-0400

Let's A, B, C had a, b, c at the start accordingly.


B loses P350 of his money to A:

"a + 350, b - 350, c;\\
a+350=2(b-350)"


Then, C loses P700 to B:

"a+350, b-350+700=b+350, c-700;\\
c-700=\cfrac{1}{3}(b+350)"


A loses P210 to C:

"a+350-210=a+140,b+350,c-700+210=c-490;\\
a+140=c-490."


So,

"\begin{cases}
 a+350=2b-700 \\
 3c-2100=b+350\\
 a+140=c-490
\end{cases}"


"\begin{cases}
 a&-2b&&=-1050\\
 &-b&+3c&=2450\\
 a&&-c&=-630
\end{cases}"


"\begin{cases}
 a-2b=-1050 \\
b=3c-2450\\
 a=c-630
\end{cases}"


"c-630-2\cdot(3c-2450)=-1050\\
-5c=-1050+630-4900\\
-5c=-5320\\c=1064\\
a=1064-630=434\\
b=3\cdot1064-2450=742."


At the start A had P434, B had P742, C had 1064.